OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = 2x2 + 6x + 4?

OpenStudy (gorv):

y=4 x=-2 and x=-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2x^2+6x+4=0 2(x+2)(x+1)=0 x=-2 or x=-1 and x=0 y=2*0^2+6*0+4=4

OpenStudy (anonymous):

could you helpme with a few more

OpenStudy (anonymous):

Solve. 2x2 – 8x – 12 = 0 x = 1 ± 2 x = 2 ± x = 2 ± x = 2 ± 2

OpenStudy (anonymous):

2(x^2-4x-6)=0 x=2+sqrt10 or x=2-sqrt10